OpenStudy (anonymous):

You can model the population of a certain city between 1955 - 2000 by the radical function P(X) = 55000 (square root x) - 1945. Using this model in which year was the population of that city 220,000? 1949 1961 1953 1957 I'm not really sure what to Assign P or X

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

if P(x) represents the population, find the x where the pop. is 220,000

OpenStudy (anonymous):

Oh ok thanks!

OpenStudy (anonymous):

200,00(x) = 55,000 (Square root x) - 1945

OpenStudy (anonymous):

woops *200,000(x) my bad

OpenStudy (bibby):

close. \(P(X) = 55000 (\sqrt x) - 1945\) \(200,000 = 55000 (\sqrt x) - 1945\)

OpenStudy (anonymous):

oh so the population is P(X) not just P okay. thank you again!

OpenStudy (anonymous):

that didn't seem to work.

OpenStudy (bibby):

P is a function. that means it takes an input,x, and has a corresponding output. P(1) is a valid statement. it means replace x with 1

OpenStudy (bibby):

it wouldn't work. they asked for x not an equation solve for x

OpenStudy (anonymous):

x = 13.482

OpenStudy (bibby):

was that wrong?

OpenStudy (anonymous):

oh wait lemme see something

OpenStudy (anonymous):

Yea it's not correct

OpenStudy (bibby):

well it's a year

OpenStudy (anonymous):

My answers are 1949 1961 1953 1957